# Quantum Corrections to $\eta/s$ from JT Gravity

### v2

@@ A.2 Quantum Green’s function for $T\ll\frac{1}{C}$
$f(M)=\log{2CM}+\pi\frac{\omega}{2}\sqrt{\frac{2C}{M}}-M\beta+{\rm ln}\Gamma\left(\ell+i\frac{\omega}{2}\sqrt{\frac{2C}{M}}\right)+{\rm ln}\Gamma\left(\ell-i\frac{\omega}{2}\sqrt{\frac{2C}{M}}\right)\,.$ (A.9)
@@
$\approx$ $e^{-\frac{2\pi^{2}C}{\beta}}\frac{1}{\pi^{3/2}}\left(\frac{\beta}{2C}\right)^{\frac{3}{2}}\frac{4}{(2C)^{2\ell-2}}\frac{(\ell-1)!^{2}}{\Gamma(2\ell)}\int d\omega e^{-i\omega t}\left(\frac{\beta}{2C}\right)^{-1}e^{\pi\frac{\omega}{2}\sqrt{2C\beta}}e^{-1}$
@@
$\approx$ $\frac{4}{(2C)^{2\ell-2}}\frac{1}{\pi^{3/2}}\frac{(\ell-1)!^{2}}{\Gamma(2\ell)}e^{-1}\int d\omega e^{-i\omega t}\left(\frac{\beta}{2C}\right)^{\frac{1}{2}}e^{\pi\frac{\omega}{2}\sqrt{2C\beta}}$
@@
$=$ $\frac{(\ell-1)!^{2}}{2^{2\ell-1}(2C)^{-1}}\frac{8}{\sqrt{\pi}e}\left(\frac{\beta}{2C}\right)^{\ell}\int\frac{d\omega}{2\pi}e^{-i\omega t}\left(\frac{2\pi}{\pi\sqrt{2C\beta}}\right)^{2\ell-1}e^{\frac{\omega}{2}\pi\sqrt{2C\beta}}$
@@
$=$ $\frac{(\ell-1)!^{2}}{2^{2\ell-1}(2C)^{-1}}\frac{8}{\sqrt{\pi}e}\left(\frac{\beta}{2C}\right)^{\ell}\left(\frac{\pi}{\pi\sqrt{2C\beta}{\rm sinh}(\frac{\pi}{\pi\sqrt{2C\beta}}t)}\right)^{2\ell}\,$, (A.12)
@@
where in the last equality we again have used the integral method from Lam:2018pvp . For the retarded Green’s function, following Faulkner:2011tm we perform a Fourier transformation. As in Section A.1, we again reinstate the constant $B$ and use the same rescaling as in Lam:2018pvp , to finally obtain
@@
$\langle\mathcal{G}_{R}(\omega,\beta)\rangle\approx-\frac{(\ell-1)!^{2}}{(2C)^{2\ell-1}}\frac{8}{\sqrt{\pi}e}\left(\frac{\beta}{2C}\right)^{\frac{1}{2}}\frac{\Gamma(\ell-i\frac{\omega}{2\pi}\pi\sqrt{2C\beta})}{\Gamma(1-\ell-i\frac{\omega}{2\pi}\pi\sqrt{2C\beta})}\,.$ (A.13)

### v3

@@ A.2 Quantum Green’s function for $T\ll\frac{1}{C}$
$f(M)=\ln{2CM}+\pi\frac{\omega}{2}\sqrt{\frac{2C}{M}}-M\beta+{\rm ln}\Gamma\left(\ell+i\frac{\omega}{2}\sqrt{\frac{2C}{M}}\right)+{\rm ln}\Gamma\left(\ell-i\frac{\omega}{2}\sqrt{\frac{2C}{M}}\right)\,.$ (A.12)
@@
$\approx$ $e^{-\frac{2\pi^{2}C}{\beta}}\frac{1}{\pi^{3/2}}\left(\frac{\beta}{2C}\right)^{\frac{3}{2}}\frac{4}{(2C)^{2\ell-2}}\frac{(\ell-1)!^{2}}{\Gamma(2\ell)}\int d\omega e^{-i\omega t}\left(\frac{\beta}{2C}\right)^{-1}\frac{\sqrt{2\pi}}{\beta}e^{\pi\frac{\omega}{2}\sqrt{2C\beta}}e^{-1}$
@@
$\approx$ $\frac{4}{(2C)^{2\ell-2}}\frac{1}{\pi^{3/2}}\frac{(\ell-1)!^{2}}{\Gamma(2\ell)}e^{-1}\int d\omega e^{-i\omega t}\left(\frac{\beta}{2C}\right)^{\frac{1}{2}}\frac{\sqrt{2\pi}}{\beta}e^{\pi\frac{\omega}{2}\sqrt{2C\beta}}$
@@
$=$ $\frac{(\ell-1)!^{2}}{2^{2\ell-9/2}e}\left(\frac{\beta}{2C}\right)^{\ell-1}\int\frac{d\omega}{2\pi}e^{-i\omega t}\left(\frac{2\pi}{\pi\sqrt{2C\beta}}\right)^{2\ell-1}e^{\frac{\omega}{2}\pi\sqrt{2C\beta}}$
@@
$=$ $\frac{(\ell-1)!^{2}}{2^{2\ell-9/2}e}\left(\frac{\beta}{2C}\right)^{\ell-1}\left(\frac{\pi}{\pi\sqrt{2C\beta}{\rm sinh}(\frac{\pi}{\pi\sqrt{2C\beta}}t)}\right)^{2\ell}\,$, (A.15)
@@
where in the last equality, we have used the integral method from Lam:2018pvp . We perform a Fourier transformation of (A.15) and obtain the Wightman function,
$\langle\mathcal{G}(\omega)\rangle$
$=\frac{(\ell-1)!^{2}}{2^{2\ell-9/2}e}\left(\frac{\beta}{2C}\right)^{\ell-1}\frac{1}{2\pi}\left(\frac{2\pi}{\pi\sqrt{2C\beta}}\right)^{2\ell-1}e^{\frac{\omega}{2}\pi\sqrt{2C\beta}}\frac{\Gamma\left(\ell+i\frac{\omega}{2\pi}\pi\sqrt{2C\beta}\right)\,\Gamma\left(\ell-i\frac{\omega}{2\pi}\pi\sqrt{2C\beta}\right)}{\Gamma(2\ell)}$
$=\frac{(\ell-1)!^{2}}{2^{2\ell-9/2}e}\frac{1}{\Gamma(2\ell)}\left(\frac{\beta}{2C}\right)^{\ell-1}\frac{1}{2\pi}\left(\frac{2\pi}{\pi\sqrt{2C\beta}}\right)^{2\ell-1}e^{\frac{\omega}{2}\pi\sqrt{2C\beta}}\frac{\pi}{{\rm sin}(\pi\ell+\pi i\frac{\omega}{2}\sqrt{2C\beta})}\frac{\Gamma\left(\ell-i\frac{\omega}{2\pi}\pi\sqrt{2C\beta}\right)}{\Gamma(1-\ell-i\frac{\omega}{2}\sqrt{2C\beta})}\,$, (A.16)
where in the last equality we used the identity
$\Gamma(1-\ell-i\frac{\omega}{2}\sqrt{2C\beta})\Gamma(\ell+i\frac{\omega}{2}\sqrt{2C\beta})=\frac{\pi}{{\rm sin}(\pi\ell+\pi i\frac{\omega}{2}\sqrt{2C\beta})}\,.$ (A.17)
To obtain the imaginary part of the retarded Green’s function $\langle\mathcal{G}_{R}(\omega,\beta)\rangle$, we use the KMS condition
@@
${\rm Im}\;\langle\mathcal{G}_{R}(\omega,\beta)\rangle=\frac{1}{2}\left(1-e^{-\beta\omega}\right)\langle\mathcal{G}(\omega)\rangle$
$=\frac{1}{2}\left(1-e^{-\beta\omega}\right)\frac{(\ell-1)!^{2}}{2^{2\ell-9/2}e}\frac{1}{\Gamma(2\ell)}\left(\frac{\beta}{2C}\right)^{\ell-1}\frac{1}{2\pi}\left(\frac{2\pi}{\pi\sqrt{2C\beta}}\right)^{2\ell-1}$
$\times e^{\frac{\omega}{2}\pi\sqrt{2C\beta}}\frac{\pi}{{\rm sin}(\pi\ell+\pi i\frac{\omega}{2}\sqrt{2C\beta})}\frac{\Gamma\left(\ell-i\frac{\omega}{2\pi}\pi\sqrt{2C\beta}\right)}{\Gamma(1-\ell-i\frac{\omega}{2}\sqrt{2C\beta})}\,.$ (A.18)
Since for positive integer $\ell$, the retarded Green’s function should be purely imaginary, one is free to write the total retarded Green’s function as
$\langle\mathcal{G}_{R}(\omega,\beta)\rangle$ $=i\frac{1}{2}\left(1-e^{-\beta\omega}\right)\frac{(\ell-1)!^{2}}{2^{2\ell-9/2}e}\frac{1}{\Gamma(2\ell)}\left(\frac{\beta}{2C}\right)^{\ell-1}\frac{1}{2\pi}\left(\frac{2\pi}{\pi\sqrt{2C\beta}}\right)^{2\ell-1}$
$\times e^{\frac{\omega}{2}\pi\sqrt{2C\beta}}\frac{\pi}{{\rm sin}(\pi\ell+\pi i\frac{\omega}{2}\sqrt{2C\beta})}\frac{\Gamma\left(\ell-i\frac{\omega}{2\pi}\pi\sqrt{2C\beta}\right)}{\Gamma(1-\ell-i\frac{\omega}{2}\sqrt{2C\beta})}\,.$ (A.19)
A.3 Modified KMS condition
From the form of the Wightman function in (A.15), we see that, after a Wick rotation, the period of the Green’s function is $\pi\sqrt{2C\beta}$. Notably, this differs from the period $\beta$ of the Green’s function in the semiclassical limit. From our calculation we see that this difference arises from a transition between different saddle points. Specifically, when moving from the semiclassical regime $(CT\gg 1)$ to the quantum regime $(CT\ll 1)$, the saddle point of the quantum-corrected Green’s function changes from $M\sim T^{2}$ to $M\sim T$. Given that the period is modified, it would be natural to define an effective inverse temperature $\beta_{\rm eff}$,
$\beta_{\rm eff}=\pi\sqrt{2C\beta}\,$, (A.20)
which could emerge due to nontrivial IR dynamics in the quantum phase.
Since the KMS condition (which relates the retarded and Wightman Green’s functions) is based on the period of the Wightman Green’s function, one might be led to adopt the following modified KMS condition,
${\rm Im}\;\langle\mathcal{G}_{R}(\omega,\beta)\rangle=\frac{1}{2}\left(1-e^{-\beta_{\rm eff}\omega}\right)\langle\mathcal{G}(\omega)\rangle\,.$ (A.21)
Using (A.21), instead of the standard prescription in terms of $\beta$, would then yield the following retarded Green’s function:
${\rm Im}\;\langle\mathcal{G}_{R}(\omega,\beta)\rangle=\frac{1}{2}\left(1-e^{-\beta_{\rm eff}\omega}\right)\langle\mathcal{G}(\omega)\rangle$
$=\frac{1}{2}\left(1-e^{-\beta_{\rm eff}\omega}\right)\frac{(\ell-1)!^{2}}{2^{2\ell-9/2}e}\frac{1}{\Gamma(2\ell)}\left(\frac{\beta}{2C}\right)^{\ell-1}\frac{1}{2\pi}\left(\frac{2\pi}{\pi\sqrt{2C\beta}}\right)^{2\ell-1}$
$\times e^{\frac{\omega}{2}\pi\sqrt{2C\beta}}\frac{\pi}{{\rm sin}(\pi\ell+\pi i\frac{\omega}{2}\sqrt{2C\beta})}\frac{\Gamma\left(\ell-i\frac{\omega}{2\pi}\pi\sqrt{2C\beta}\right)}{\Gamma(1-\ell-i\frac{\omega}{2}\sqrt{2C\beta})}\,.$ (A.22)
Taking $\ell=1$ and working in the $\omega\ll T\ll C^{-1}$ limit, the retarded Green’s function above can be approximated as follows,
${\rm Im}\langle\mathcal{G}_{R}(\omega,\beta)\rangle$ $=\frac{2^{3/2}}{e}e^{\frac{\omega}{2}\pi\sqrt{2C\beta}}\frac{\pi}{{\rm sinh}(\pi\frac{\omega}{2}\sqrt{2C\beta})}\left(\frac{\omega}{2\pi}\right)\left(1-e^{-\beta_{\rm eff}\omega}\right)$
$\approx\frac{2^{3/2}}{e}\omega\,.$ (A.23)
Finally, the corresponding shear viscosity in the quantum regime $T\ll\frac{1}{C}$ would be
$\eta^{\prime}=\frac{1}{2\kappa^{2}}\left(\frac{r_{h}}{L}\right)^{2}\times\frac{2^{3/2}}{e}\,$, (A.24)
which is a constant and therefore doesn’t exhibit temperature dependence in the quantum regime. We note that this result would not match the absorption cross sections of Emparan:2025sao ; Biggs:2025nzs , which were derived assuming the standard KMS condition.
It would be interesting to have a better understanding [^14] We are grateful to Elias Kiritsis, Hong Liu and Sameer Murthy for comments on the KMS condition in this case. of the physical origin of the effective inverse temperature $\beta_{\rm eff}$. This may imply the existence of a transition in the dominant spacetime geometry, from the semi-classical regime to the fully quantum regime [^15] Similarly, to resolve the negative entropy issue in the quantum limit, one should take into account the contribution of replica wormholes; see Chandrasekaran:2022asa .. It may also arise from the approximations used to obtain (A.15), which might not reflect features visible in the full integral. However, due to the limited understanding of hydrodynamic calculations in the quantum regime, and in order to reproduce the previous results for the cross section, in the main text we have adopted the standard KMS prescription with $\beta$. We look forward to having a clearer interpretation of $\beta_{\mathrm{eff}}$.